### v1

@@ Abstract
Given a sequence of tokens, such as words, the task of next-token prediction is to predict the next-token conditional probability distribution. Decoder-only transformers have become effective models for this task, but their properties are still not fully understood. In particular, the largest number of distinct context sequences that a decoder-only transformer can interpolate next-token distributions for has not been established. To fill this gap, we prove upper and lower bounds on this number, which are equal up to a multiplicative constant. We prove these bounds in the general setting where next-token distributions can be arbitrary as well as the empirical setting where they are calculated from a finite number of document sequences. Our lower bounds are for one-layer transformers and our proofs highlight an important injectivity property satisfied by self-attention. Furthermore, we provide numerical evidence that the minimal number of parameters for memorization is sufficient for being able to train the model to the entropy lower bound.
@@ 1 Introduction
Ten years after Bengio et al (2000) introduced neural probabilistic language models, Tomáš Mikolov constructed a language model that takes a sequence of tokens (e.g. words) as input, embeds the tokens into a vector space using a Skip-gram model Mikolov et al (2013), then inputs the sequence of token vectors to a recurrent neural network (RNN) (Mikolov, 2012). Three years later, the attention mechanism was developed for combining a sequence of vectors into a single vector (Bahdanau et al, 2015). One year later, self-attention was developed as a modification to attention (Cheng et al, 2016). One year after that, Vaswani et al (2017) recognized that self-attention made it unnecessary to use a RNN and so replaced the RNN with a feedforward neural network (FNN), introducing the modern transformer model. This gives a brief timeline of the progress leading to modern transformers. But language models have been around for much longer.
@@
75 years ago, Claude Shannon introduced the concept of the entropy of language and used the predictability of language to study it (Shannon, 1948, 1951). The entropy of a discrete probability distribution, $P$, over $\omega$ possible events, is the quantity $-\sum_{\gamma=1}^{\omega}P(\gamma)\log(P(\gamma))$, and it satisfies certain properties such that it intuitively captures “how much ‘choice’ is involved in the selection of the event or of how uncertain we are of the outcome” (Shannon, 1948). If we are trying to fit a probability distribution, $Q_{\theta}$, parameterized by the vector $\theta\in\mathbb{R}^{k}$, to $P$, then we may use the cross-entropy loss, $-\sum_{\gamma=1}^{\omega}P(\gamma)\log(Q_{\theta}(\gamma))$, which is greater than or equal to the entropy of $P$ by Gibb’s inequality. More generally, the entropy of a probability distribution, $P$, over sequences of $\omega$ possible events, is $-\sum_{t=1}^{\infty}\sum_{\alpha\in[\omega]^{t}}\sum_{\gamma=1}^{\omega}P(\alpha,\gamma)\log(P(\alpha,\gamma\mid\alpha))$, and it lower bounds the cross-entropy $-\sum_{t=1}^{\infty}\sum_{\alpha\in[\omega]^{t}}\sum_{\gamma=1}^{\omega}P(\alpha,\gamma)\log(Q_{\theta}(\alpha,\gamma\mid\alpha))$. The idea of “next-token prediction” is to tune $\theta\in\mathbb{R}^{k}$ to minimize the cross-entropy.
Of course, in practice, $P$ is replaced with the empirical next-token distribution $\hat{P}$. For example, given a data set of $N$ document sequences $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$, the cross-entropy loss is proportional to
@@
$-\sum_{i=1}^{N}\sum_{t=1}^{T}\log(Q_{\theta}(\beta_{1}^{i},\ldots,\beta_{t+1}^{i}\mid\beta_{1}^{i},\ldots,\beta_{t}^{i}))$. If $\hat{P}(\alpha)$ is nonzero for $n$ $\alpha$’s in $\cup_{t=1}^{T}[\omega]^{t}$, i.e. if the document sequences have $n$ unique contexts, $\alpha^{i}=(\beta_{1}^{j},\ldots,\beta_{t}^{j})$, then we show, in Theorem 1, that $\approx n\omega$ parameters is necessary for the infimum of the cross-entropy loss to equal the entropy of $\hat{P}$. Then, in Theorem 5, we show that when $Q_{\theta}$ is a one-layer decoder-only transformer, $\approx n\omega$ parameters is actually sufficient. Interestingly, this requires that the number of feedforward neural-network (FNN) neurons be greater than or equal to $n$, which means that the FNN is able to smoothly interpolate the output vectors from the self-attention sub-layer (Bubeck and Sellke, 2021).
@@
A one-layer decoder-only transformer model has three sub-layers: the embedding, the (multi-head) self-attention, and the FNN. We will show that this model has optimal interpolation power by first showing that it already does with embedding dimension 1 and only one self-attention head. In this setting, the token embeddings and positional embeddings are scalars. Given an input sequence, the self-attention sub-layer outputs a scalar as well. In Theorem 4, we show that it does so injectively, which allows the model to exploit the interpolation power of the FNN. This leads us to consider simple alternatives to self-attention that satisfy the same injectivity property. As it turns out, taking the weighted average of embedded token sequences, which we call “token averaging,” works as well, as we show in Theorem 6. Thus, replacing self-attention with token averaging in the one-layer decoder-only transformer model, results in another model with optimal memory capacity: Theorem 7.
@@
The memory capacity, the largest $n$ such that a model can interpolate $n$ generic input-output pairs, is a classic topic in learning theory (Cover, 1965), but it is only recently, with the shift in perspective from the classical bias-variance trade-off to the double descent curve of Belkin et al (2019), that going beyond the interpolation threshold has been viewed as a benefit to generalization. Belkin et al (2019) argue that when there is an abundance of interpolating solutions, simple algorithms such as gradient descent tend to choose the smoothest interpolating solution, which is a form of Occam’s razor and therefore may be at the heart of the ability for highly expressive models to generalize beyond the data sets they are trained on. Thus, it is even reasonable to construct a model by thinking about the simplest way to achieve optimal memory capacity, so it is especially important to investigate the memory capacity of models which are already used in practice.
@@ 1.1 Results
We prove upper and lower bounds on the memory capacity of one-layer decoder-only transformers for next-token prediction in two settings: the general setting where the next-token distribution conditioned on a context sequence can be arbitrary, and the empirical setting where it is calculated from a finite number of document sequences. Let $k$ be the number of model parameters and $n$ be the number of distinct context sequences. Then our upper and lower bounds are summarized in Table 1. The particular constant hidden by $\Omega(\cdot)$ is given in Equation (3).
@@
Memory capacity bounds for transformers
@@
Upper bounds $\frac{k}{\omega-1}$, Lemma 1 $\left(2+\frac{1}{\omega-1}\right)\frac{k}{\omega-1}+1$, Theorem 1
Lower bounds $\Omega\left(\frac{k}{\omega-1}\right)$, Theorem 5 $\Omega\left(\frac{k}{\omega-1}\right)$, Lemma 2 applied to Theorem 5
We prove the memory capacity lower bound in Theorem 5 by reducing to the one-head setting with embedding dimension 1. In this setting, the self-attention sub-layer outputs scalars to the FNN sub-layer. Given $n$ distinct context sequences, let $x\in\mathbb{R}^{n}$ denote the $n$ scalar outputs. Then the model has the form $V^{\top}\psi(wx^{\top})$ where $\psi$ is the activation function of the FNN. If $\psi(wx^{\top})$ has rank $n$, then we can solve for $V$ given any target $Y\in\mathbb{R}^{\omega\times n}$. Thus, the memory capacity of $V^{\top}\psi(wx^{\top})$ comes down to the rank of $\psi(wx^{\top})$.
In Theorem 2, we show that if $\psi$ is a polynomial with $k$ monomials and $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct, then the rank and Kruskal rank of $\psi(ab^{\top})$ are $\min\{m,n,k\}$ for generic $a\in\mathbb{R}^{m}$. In Theorem 3, we show that if $\psi$ is real analytic at zero and not a polynomial there, and $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct, then the rank and Kruskal rank of $\psi(ab^{\top})$ are $\min\{m,n\}$ for generic $a$ in a neighborhood of zero in $\mathbb{R}^{m}$.
Thus, the memory capacity of $V^{\top}\psi(wx^{\top})$ comes down to whether or not the entries of $x$ are nonzero and distinct. To show that they are, we show, in Theorem 4, that self-attention maps injectively into \$\0\mathbb{R}\backslash\{0\}$. This also leads us to consider token averaging as a simple alternative to self-attention. In Theorem 6, we show that token averaging also maps injectively into \$\0\mathbb{R}\backslash\{0\}$, which leads to an optimal memory capacity lower bound, Theorem 7, for the composition of token averaging and a FNN.
@@ 1.2 Related work
The memory capacity of a machine learning model with $k$ parameters is the largest $n$ such that it can interpolate $n$ generic input-output pairs (Cover, 1965). Baum (1988) showed that a two-layer neural network with Heaviside activation has memory capacity at least $\approx k$ where outputs are in {±$\{\pm 1\}$. Yun et al (2019) proved a similar result but with rectified linear unit (ReLU) activation. Then, a year later, Bubeck et al (2020) showed that this actually applies more generally for outputs in $\mathbb{R}$. Here, because $\mathbb{R}^{k}$ and $\mathbb{R}^{n}$ are both manifolds, $k\geq n$ is necessary to interpolate $n$ generic input-output pairs, and so the memory capacity result is optimal. Four years later, Madden and Thrampoulidis (2024) proved a similar result but for arbitrary activation functions that are real analytic at a point and, if reducing to a polynomial there, have sufficiently high degree. Thus, the optimal memory capacity of two-layer neural networks has been well established.
@@
In the setting of next-token prediction, inputs are context sequences in $\cup_{t=1}^{\infty}[\omega]^{t}$ and outputs are in $\Delta^{\omega-1}$, the unit simplex in $\mathbb{R}^{\omega}$. An output is viewed as a prediction of the distribution of the next-token conditioned on the context sequence. While Kim et al (2023), Mahdavi et al (2024), and Kajitsuka and Sato (2024) prove memory capacity results about transformers, only Mahdavi et al (2024) views inputs as context sequences. Kim et al (2023) and Kajitsuka and Sato (2024) prove memory capacity results in the sequence-to-sequence setting, where inputs are document sequences of token embeddings in $\mathbb{R}^{d\times T}$ and outputs are in $[C]^{1\times T}$ or $[C]^{d\times T}$ for some $C\in\mathbb{N}$. This does not apply to the next-token prediction setting for two reasons. First, the number of document sequences, $N$, could go to infinity while the number of unique contexts, $n$, remains fixed, thus the number of parameters required to apply the result, $\Omega(\sqrt{NT})$ in Kim et al (2023) and $\Omega(NT)$ in Kajitsuka and Sato (2024), is vacuous in the next-token prediction setting. Second, the outputs are not in $\mathbb{R}^{\omega\times T}$. Remark 3.2 of Kim et al (2023) explains how the codomain can be extended to ×$\mathbb{R}^{\ell\times T}$ for any $\ell\in\mathbb{N}$ by fixing the precision level, $\epsilon>0$, but the new required number of parameters is $\Omega(\ell\log(\ell/\epsilon)\sqrt{NT})$, which goes to infinity as $\epsilon$ goes to zero. Furthermore, both of these results assume that the embedded tokens are sufficiently separated while we do not make any assumptions on inputs.
@@
Thus, the result of Mahdavi et al (2024) is the most relevant to ours. They prove a memory capacity result for transformers in the setting where inputs are context sequences of token embeddings in $\mathbb{R}^{d\times\tau}$ and each output is a prediction of the next-token embedding in $\mathbb{R}^{d}$, showing that $\approx nd$ parameters are sufficient to interpolate $n$ input-output pairs, which is optimal. Interestingly, they analyzed a model without a FNN sub-layer and instead leveraged the number of self-attention heads to prove their result. However, besides only applying to outputs in $\mathbb{R}^{d}$, rather than $\mathbb{R}^{\omega}$, their result requires that all context sequences have the same length.
In conclusion, while significant progress has been made in the analysis of the memory capacity of transformers in just the past year, our Theorems 5 is the first optimal memory capacity lower bound for transformers in the next-token prediction setting. Our upper bound, Theorem 1, can be seen as a generalization of Theorem 1 of Sontag (1997) from input-output pairs in $\mathbb{R}^{d}\times\mathbb{R}$ to input-output pairs in $\cup_{t=1}^{\infty}[\omega]^{t}\times\Delta^{\omega-1}$. However, our proof is quite different because $\cup_{t=1}^{\infty}[\omega]^{t}$ is countable.
As for Theorems 2 and 3, these are classical results when $a$ and $b$ are both generic, but to the best of our knowledge, they are novel results in the setting where $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct. However, closely related results have been shown before. Tamura (1991) proved a version of Theorem 3 when $\psi$ is the logistic function and restated the result in Tamura and Tateishi (1997). Then Huang and Babri (1998) proved it more generally when $\psi$ is bounded, nonlinear, and has a limit at either positive or negative infinity. Their result does not require $\psi$ to be real analytic at a point and not a polynomial there, but our result does not require $\psi$ to be bounded or to have limits. Thus, the two results are complementary. Interestingly, their proof sends the input values of $\psi$ off to infinity (or negative infinity), while our proof only uses input values in the interval of convergence of the point where $\psi$ is real analytic.
@@
We have reviewed the works most closely related to our memory capacity results, but there are many relevant studies exploring the optimization of one-layer transformers as well, such as Tian et al (2023); Chen and Li (2024); Tarzanagh et al (2023); Zhang et al (2023). However, most do not address the next-token prediction scenario we study here. The exceptions, such as Makkuva et al (2024); Li et al (2024); Tian et al (2023, 2024); Ildiz et al (2024), rely on restrictive assumptions about input data and transformer dimensions, such as embedding size, to achieve minimal empirical risk. Additionally, optimization and learnability in the next-token prediction setting, albeit limited to linear models, have been recently studied in Malach (2023); Thrampoulidis (2024).

### v2

@@ Abstract
Given a sequence of tokens, such as words, the task of next-token prediction is to predict the next-token conditional probability distribution. Decoder-only transformers have become effective models for this task, but their properties are still not fully understood. In particular, the largest number of distinct context sequences that a decoder-only transformer can interpolate next-token distributions for has not been established. To fill this gap, we prove upper and lower bounds on this number, which are equal up to a multiplicative constant. We prove these bounds in the general setting where next-token distributions can be arbitrary as well as the empirical setting where they are calculated from a finite number of document sequences. Our lower bounds are for one-layer multi-head decoder-only transformers and our proofs highlight an important injectivity property satisfied by self-attention. Furthermore, we provide numerical evidence that the minimal number of parameters for memorization is sufficient for being able to train the model to the entropy lower bound.
@@ 1 Introduction
Ten years after Bengio et al (2000) introduced neural probabilistic language models, Tomáš Mikolov constructed a language model that takes a sequence of tokens (e.g. words) as input, embeds the tokens into a vector space using a Skip-gram model, then inputs the sequence of token vectors to a recurrent neural network (RNN) (Mikolov, 2012). Three years later, the attention mechanism was developed for combining a sequence of vectors into a single vector (Bahdanau et al, 2015). One year later, self-attention was developed as a modification to attention (Cheng et al, 2016). One year after that, Vaswani et al (2017) recognized that self-attention made it unnecessary to use a RNN and so replaced the RNN with a feedforward neural network (FNN), introducing the modern transformer model. This gives a brief timeline of the progress leading to modern transformers. But language models have been around for much longer.
@@
75 years ago, Claude Shannon introduced the concept of the entropy of language and used the predictability of language to study it (Shannon, 1948, 1951). The entropy of a probability mass function $p$ over a finite set $\mathcal{X}$ is the quantity $-\sum_{x\in\mathcal{X}}p(x)\log p(x)$ and it satisfies certain properties such that it intuitively captures “how much ‘choice’ is involved in the selection of the event or of how uncertain we are of the outcome” (Shannon, 1948). If we are trying to fit a probability mass function, $q_{\theta}$, parameterized by the vector $\theta\in\mathbb{R}^{k}$, to $p$, then we can do so by minimizing the cross-entropy of $q_{\theta}$ relative to $p$, $-\sum_{x\in\mathcal{X}}p(x)\log q_{\theta}(x)$, which is greater than or equal to the entropy of $p$ by Gibb’s inequality. In the setting of next-token prediction, we have a vocabulary, $[\omega]=\{1,\ldots,\omega\}$, and we assume that each context sequence $\alpha\in\cup_{t=0}^{T-1}[\omega]^{t}$ has a corresponding next-token probability mass function $p(\cdot\mid\alpha)$ over $[\omega]$. Here, the cross-entropy of $q_{\theta}$ relative to $p$ is defined as $-\sum_{t=1}^{T}\sum_{\alpha\in[\omega]^{t-1}}\sum_{\gamma=1}^{\omega}p(\alpha,\gamma)\log q_{\theta}(\gamma\mid\alpha)$.
@@
However, in practice, instead of being given $p$, we are generally given a text corpus of $N$ document sequences $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T}$. In this case, it turns out that the cross-entropy of $q_{\theta}$ relative to the empirical $\hat{p}$, calculated from the text corpus, is equal to $-\frac{1}{N}\sum_{t=1}^{T}\sum_{j=1}^{N}\log q_{\theta}(\beta_{t}^{j}\mid\beta_{1}^{j},\ldots,\beta_{t-1}^{j})$, and so $\hat{p}$ does not have to be computed. If $\hat{p}(\alpha)$ is nonzero for $n$ $\alpha$’s in $\cup_{t=1}^{T-1}[\omega]^{t}$, then we show, in Theorem 4.5, that $\approx n\omega$ parameters is necessary for the infimum of the cross-entropy loss to equal the entropy of $\hat{p}$. Then, in Theorem 6.5, we show that when $q_{\theta}$ is a one-layer multi-head decoder-only transformer, $\approx n\omega$ parameters is actually sufficient.
@@
A one-layer multi-head decoder-only transformer model has three sub-layers: the embedding, the (multi-head) self-attention, and the FNN. We will show that this model has optimal interpolation power by first showing that it already does with embedding dimension 1 and only one self-attention head. In this setting, the token embeddings and positional embeddings are scalars. Given an input sequence, the self-attention sub-layer outputs a scalar as well. In Theorem 6.1, we show that it does so injectively, which allows the model to exploit the interpolation power of the FNN. This leads us to consider simple alternatives to self-attention that satisfy the same injectivity property. As it turns out, taking the weighted average of embedded token sequences, which we call token-averaging, works as well, as we show in Theorem 8.1. Thus, replacing self-attention with token-averaging in the transformer model results in another model with optimal memory capacity, as we show in Theorem 8.2.
@@
The memory capacity, the largest $n$ such that a model can interpolate $n$ generic input-output pairs, is a classic topic in learning theory (Cover, 1965), but it is only recently, with the shift in perspective from the classical bias-variance trade-off to the double descent curve of Belkin et al (2019), that going beyond the interpolation threshold has been viewed as a benefit to generalization. Belkin et al (2019) argue that when there is an abundance of interpolating solutions, simple algorithms such as gradient descent tend to choose the smoothest interpolating solution, which is a form of Occam’s razor and therefore may be at the heart of the ability for highly expressive models to generalize beyond the data sets they are trained on.
@@ 1.1 Results
We formalize the idea of a probabilistic language space and show that next-token distributions are sufficient to fully specify an underlying stochastic process for language. In particular, Propositions 3.4 and 3.5 prove that the extension from next-token distributions to a probability measure on the cylindrical $\sigma$-algebra is unique by: (1) showing that next-token distributions correspond to probabilities of thin cylinders, (2) showing that the system of the empty set and thin cylinders is a $\pi$-system, and (3) applying Dynkin’s $\pi$-$\lambda$ theorem.
Then we formally define next-token prediction capacity as the memory capacity of a next-token prediction model. We also formally define empirical next-token prediction capacity which applies to the setting where the data comes from a text corpus. We upper bound the next-token prediction capacity for continuously differentiable models in Lemma 4.2, upper bound the next-token prediction capacity by the empirical next-token prediction capacity for any model in Lemma 4.4, and upper bound the empirical next-token prediction capacity for continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable models in Theorem 4.5.
Then we focus our attention on the one-layer multi-head decoder-only transformer model. We lower bound the next-token prediction capacity for this model in Theorem 6.5. Let $k$ be the number of model parameters and $n$ be the number of distinct context sequences. Then our upper and lower bounds are summarized in Table 1. The particular constant hidden by $\Omega(\cdot)$ is given in Equation (7).
@@
Next-token prediction capacity bounds for a transformer with $k$ parameters
@@
Upper bounds $\frac{k}{\omega-1}$, Lemma 4.2 $\left(2+\frac{1}{\omega-1}\right)\frac{k}{\omega-1}+1$, Theorem 4.5
Lower bounds $\Omega\left(\frac{k}{\omega-1}\right)$, Theorem 6.5 $\Omega\left(\frac{k}{\omega-1}\right)$, Lemma 4.4 + Theorem 6.5
We prove the next-token prediction capacity lower bound in Theorem 6.5 by reducing to the setting with embedding dimension equal to one. In this setting, the self-attention sub-layer outputs scalars to the FNN sub-layer. Given $n$ distinct context sequences, let $x\in\mathbb{R}^{n}$ denote the $n$ scalar outputs. Then the model has the form $V^{\top}\psi(wx^{\top})$ where $\psi$ is the activation function of the FNN. If $\psi(wx^{\top})$ has rank $n$, then we can solve for $V$ given any target $Y\in\mathbb{R}^{\omega\times n}$. Thus, the memory capacity of $V^{\top}\psi(wx^{\top})$ comes down to the rank of $\psi(wx^{\top})$.
In Theorem 6.3, we show that if $\psi$ is a polynomial with $k$ monomials and $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct, then the rank and Kruskal rank of $\psi(ab^{\top})$ are $\min\{m,n,k\}$ for generic $a\in\mathbb{R}^{m}$. In Theorem 6.4, we show that if $\psi$ is real analytic at zero and not a polynomial there, and $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct, then the rank and Kruskal rank of $\psi(ab^{\top})$ are $\min\{m,n\}$ for generic $a$ in a neighborhood of zero in $\mathbb{R}^{m}$.
Thus, the rank of $\psi(wx^{\top})$ comes down to whether or not the entries of $x$ are nonzero and distinct. To show that they are, we prove, in Theorem 6.1, that self-attention maps injectively into \$\0\mathbb{R}\backslash\{0\}$. This also leads us to consider token-averaging as a simple alternative to self-attention. In Theorem 8.1, we show that token-averaging also maps injectively into \$\0\mathbb{R}\backslash\{0\}$, which leads to an optimal memory capacity lower bound, Theorem 8.2, for the composition of token-averaging and a FNN.
@@ 1.2 Related work
Probability of language.
One way to think about language probabilistically is to assume that for every sequence of $n-1$ tokens (whether letters or words), there is a corresponding next-token distribution. This is equivalent to assuming that there is an underlying Markov process of order $n-1$. This concept is called the $n$-gram model of language and was introduced by Markov in 1913 when he computed the $n$-gram probabilities of sequences of vowels and consonants in Pushkin’s Eugene Onegin (Markov, 2006). 35 years later, Shannon analyzed the statistical entropy of the $n$-gram model in two seminal papers, establishing the field of information theory (Shannon, 1948, 1951). The concept of statistical entropy goes back to the 1870s when Boltzmann showed that the statistical entropy of microstates in a thermodynamic system is proportional to the thermodynamic entropy of the system.
A more realistic assumption on language is that for every sequence of any length there is a corresponding next-token distribution. This idea was introduced in Shannon’s 1951 paper when he defined the entropy rate as the limit of the normalized $n$-gram entropies. From the next-token distributions, we can define a function $p$ on sequences of any length using the chain rule of probability, and this $p$ will be a probability mass function when restricted to sequences of any fixed length. We can also define a probability measure on the cylindrical $\sigma$-algebra, and from this probability measure define a stochastic process such that its finite-dimensional distributions are given by $p$. But, unlike with the $n$-gram model, it is not clear that the stochastic process is fully specified by the next-token distributions. To be precise, while the extension of $p$ to a probability measure on the cylindrical $\sigma$-algebra is standard, see e.g. the sequence space subsection of Section 2 in Billingsley (1995), the extension proceeds in two steps, first to the cylindrical algebra, then to the cylindrical $\sigma$-algebra, and only the uniqueness of the second step appears to have been noticed, rather than the uniqueness of the full extension.
Another way to think about language probabilistically is to assume that for every sequence of any length there is a corresponding probability, i.e. there is a probability mass function over sequences of any length. Booth and Thompson (1973) call such a language a probability measure language and call its probability mass function a probabilistic word function. But the sequence probabilities cannot come from next-token distributions unless a special end-of-sequence token is included to ensure that the underlying stochastic process terminates with probability one. Keeping the end-of-sequence token but relaxing the assumption that it appears in every sequence with probability one, Du et al (2023) use the extension from the previous paragraph to induce a probability measure over finite and infinite sequences simultaneously. This is an interesting direction, but is conceptually distinct from the perspective taken in the present paper.
Memory capacity.
The memory capacity—also known as finite sample expressivity, memorization capacity, storage capacity, or, simply, capacity—of a machine learning model with $k$ parameters is the largest $n$ such that it can interpolate $n$ generic input-output pairs, where by generic we mean that the set of exceptions lies on the zero set of a nontrivial real analytic function and is, therefore, measure zero and closed (Gunning and Rossi, 1965, Corollary 10). The idea of memory capacity goes back to Cover (1965) who considered the separating capacities of families of surfaces. Later, Baum (1988) proved that a two-layer FNN with Heaviside activation has memory capacity at least $\approx k$ where outputs are in {±$\{\pm 1\}$. Sakurai (1992) extended this to three-layer FNNs. Huang and Huang (1991) proved it for two-layer FNNs with sigmoid activation and outputs in $\mathbb{R}$. Yamasaki (1993) sketched a proof for $L$-layer FNNs with sigmoid activation. Huang (2003) proved it for three-layer FNNs with sigmoid activation. Yun et al (2019) proved it for two-layer FNNs with ReLU activation and outputs in {±$\{\pm 1\}$ (their Corollary 4.2), and for three-layer FNNs with ReLU activation and outputs in $\mathbb{R}$ (their Theorem 3.1). Bubeck et al (2020) proved it for two-layer FNNs with ReLU activation and outputs in $\mathbb{R}$. Madden and Thrampoulidis (2024) proved it for two-layer FNNs with general activations (excluding only low degree polynomials and low degree splines) and outputs in $\mathbb{R}$. Madden (2024) proved it for three-layer FNNs with general activations (excluding only polynomials and splines) and outputs in $\mathbb{R}$.
While all of the papers mentioned so far have been about the memory capacity of FNNs, Kim et al (2023), Kajitsuka and Sato (2024), and Mahdavi et al (2024) recently proved memory capacity lower bounds for transformers. However, none of these results apply to the next-token prediction setting where inputs are context sequences in $\cup_{t=1}^{\infty}[\omega]^{t}$ and outputs are next-token distributions in $\Delta^{\omega-1}$, the unit simplex in $\mathbb{R}^{\omega}$. We will show that this is the case for each of the three papers, but first let us clarify two points about the inputs and outputs. Regarding inputs, if $\beta=[\beta_{1}|\cdots|\beta_{T}]\in\mathbb{R}^{d\times T}$, then $\beta_{1},[\beta_{1}|\beta_{2}],[\beta_{1}|\beta_{2}|\beta_{3}],\ldots$ are its unique contexts. More generally, if $\beta^{1},\ldots,\beta^{N}\in\mathbb{R}^{d\times T}$, then $\{[\beta_{1}^{j}|\cdots|\beta_{t}^{j}]\mid j\in[N],t<T\}$ are the unique contexts of the list. If we use $n$ to denote the number of unique contexts, then it is possible for $N$ to go to infinity while $n$ remains fixed. Thus, the memory capacity for next-token prediction should be in terms of $n$, not $N$. Moving on to outputs, context sequences do not have unique next-tokens, they have next-token distributions, e.g. “I like” may be followed by “dogs” or “cats.” In other words, the next-token distributions are not necessarily one-hot. Thus, for next-token prediction, outputs should be in $\mathbb{R}^{\omega}$, not $\mathbb{R}^{d}$. Given these two subtleties, we use next-token prediction capacity to refer to the memory capacity of next-token prediction as opposed to the other settings considered in Kim et al (2023), Kajitsuka and Sato (2024), and Mahdavi et al (2024).
@@
Kim et al (2023) and Kajitsuka and Sato (2024) proved memory capacity results in the sequence-to-sequence setting, where inputs are document sequences of token embeddings in $\mathbb{R}^{d\times T}$ and outputs are in $[C]^{1\times T}$ or $[C]^{d\times T}$ for some $C\in\mathbb{N}$. This does not apply to the next-token prediction setting for two reasons. First, the number of document sequences, $N$, could go to infinity while the number of unique contexts, $n$, remains fixed, thus the number of parameters required to apply the result, $\Omega(\sqrt{NT})$ in Kim et al (2023) and $\Omega(NT)$ in Kajitsuka and Sato (2024), is vacuous in the next-token prediction setting. Second, the outputs are not in $\mathbb{R}^{\omega\times T}$. Remark 3.2 of Kim et al (2023) explains how the codomain can be extended to ×$\mathbb{R}^{\ell\times T}$ for any $\ell\in\mathbb{N}$ by fixing the precision level, $\epsilon>0$, but the new required number of parameters is $\Omega(\ell\log(\ell/\epsilon)\sqrt{NT})$, which goes to infinity as $\epsilon$ goes to zero. Furthermore, both of these results assume that the embedded tokens are sufficiently separated while we do not make any assumptions on inputs.
@@
Mahdavi et al (2024) proved a memory capacity result for transformers in the setting where inputs are context sequences of token embeddings in $\mathbb{R}^{d\times\tau}$ and each output is a prediction of the next-token embedding in $\mathbb{R}^{d}$, showing that $\approx nd$ parameters are sufficient to interpolate $n$ input-output pairs, which is optimal. Interestingly, they analyzed a model without a FNN sub-layer and instead leveraged the number of self-attention heads to prove their result. However, besides only applying to outputs in $\mathbb{R}^{d}$, rather than $\mathbb{R}^{\omega}$, their result requires that all context sequences have the same length. Thus, their result does not apply to the unique contexts $\beta_{1},[\beta_{1}|\beta_{2}],[\beta_{1}|\beta_{2}|\beta_{3}],\ldots$ of a document sequence $\beta\in\mathbb{R}^{d\times T}$, only to contexts $\alpha\in\mathbb{R}^{d\times\tau}$ for some fixed $\tau$.
Optimization of transformers.
@@
We have reviewed the works most closely related to our memory capacity results, but there are many relevant studies exploring the optimization of one-layer transformers as well, such as Tian et al (2023), Chen and Li (2024), Tarzanagh et al (2023), Zhang et al (2024), and Deora et al (2024). However, most do not address the next-token prediction scenario we study here. The exceptions, such as Makkuva et al (2024), Li et al (2024), Tian et al (2023), Tian et al (2024), and Ildiz et al (2024), rely on restrictive assumptions about input data and transformer dimensions, such as embedding size, to achieve minimal empirical risk. Additionally, optimization and learnability in the next-token prediction setting, albeit limited to linear models, have been recently studied in Malach (2023) and Thrampoulidis (2024). Finally, the effectiveness of transformers for particular tasks has been explored in papers such as Sanford et al (2023) and Wang et al (2024).
Miscellaneous results.
Our upper bound, Theorem 4.5, can be seen as a generalization of Theorem 1 of Sontag (1997) from input-output pairs in $\mathbb{R}^{d}\times\mathbb{R}$ to input-output pairs in $\cup_{t=1}^{\infty}[\omega]^{t}\times\Delta^{\omega-1}$. However, our proof is quite different because $\cup_{t=1}^{\infty}[\omega]^{t}$ is countable. As for Theorems 6.3 and 6.4, these are classical results when $a$ and $b$ are both generic, but to the best of our knowledge, they are novel results in the setting where $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct. However, closely related results have been shown before. Tamura (1991) proved a version of Theorem 6.4 when $\psi$ is the logistic function and restated the result in Tamura and Tateishi (1997). Then Huang and Babri (1998) proved it more generally when $\psi$ is bounded, nonlinear, and has a limit at either positive or negative infinity. Their result does not require $\psi$ to be real analytic at a point and not a polynomial there, but our result does not require $\psi$ to be bounded or to have limits. Thus, the two results are complementary. Interestingly, their proof sends the inputs of $\psi$ off to positive or negative infinity, while our proof only uses inputs in the interval of convergence of the point where $\psi$ is real analytic.